In a circle with 5 cm radius, an 11 cm arc subtends a central angle of how many radians? Of how many degrees, to the nearest degree?
Answer by stanbon(75887)

In a circle with 5 cm radius, an 11 cm arc subtends a central angle of how many radians? Of how many degrees, to the nearest degree?
------
arc length = (theta)r where theta is a central angle in radians and
r = radius.
------------------------
Solve:
11 cm = theta(5)
theta = 11/5 radians
----------------------------
Converting to degrees:
(11/5)(180/pi) = (11*180/(5pi)) = 126.051 degrees
================================
Cheers,
Stan H.